# Obfuscated Location Disclosure for Remote ID Enabled Drones

### v1

@@ 9 Discussion
Summary. Through our extensive performance assessment (Sec. 8), we experimentally proved that OLO-RID allows to generate and deliver RID messages protecting the drones’ location privacy within $0.16$ s, i.e., below the threshold of $1$ s set by the RID regulation [4]. The RAM consumption and communication overhead of our solution are also negligible, guaranteeing the potential smooth integration of our solution into running RID-compliant deployments.
Energy Considerations. As discussed in Sec. 8.2, a single RID message requires $0.149$ mW and $0.216$ mW of power for the ECs BN254 and BLS48556 on the RPI, respectively. Therefore, a single RID message requires $0.0140$ mJ and $0.0147$ mJ of energy for ECs BN254 and BLS48556, respectively. We can express such values in ampere-hours as $Ah=E/(V\cdot 3600)$, where $Ah$ is the ampere-hours (Ah), $E$ is the energy usage (J), and V is the voltage (V) of the battery of the UAV, e.g., $7.7$ V for the DJI Mini 2 [37]. Therefore, a single RID message secured through our solution requires from the UAV $0.000505$ $\mu$Ah and $0.000530$ $\mu$Ah for ECs BN254 and BLS48556, respectively. As the battery capacity of the DJI Mini 2 is $2250$ mAH, OLO-RID consumes between $0.0224\%$ and $0.0236\%$ of the DJI Mini 2 total battery, confirming its limited energy overhead.
@@
Compliance to Requirements. Here, we look back to the requirements defined in Sec.3.4 to evaluate the compliance of our solution. The requirement R1, Location Obfuscation is fulfilled by integrating and extending the solution in [9]. As per the requirement R2, Trajectory Privacy, as shown in Sec. 6, our system is as secure as the PIM method, so it offers trajectory privacy. Our solution does not require any Internet connection on the UAV, so it also fulfils the requirement R3, No Persistent Internet Connection. Internet connection might be required for some use cases discussed in Sec. 3.3, but the absence of such a connection does not prevent the drone from running OLO-RID. As per the requirement R4, Maximum Message Generation Time., as shown in Sec. 8.2, the maximum time our solution takes to generate a RID message is always less than $0.16$ seconds. Finally, as per the requirement R5, Maximum Communication Overhead, the modified RID message disclosed through OLO-RID is at most 227 bytes, below the MTU of WiFi of $2,304$  bytes [21].
@@
RID Extension and Backward Compatibility. To protect the UAVs’ location privacy, OLO-RID extends the format of standard RID messages. Instead of delivering the actual UAV location, we deliver an obfuscated location, i.e., a location loosely correlated with the actual one, as described in Sec. 5. We also extend the standard RID message with encrypted location reports, containing the encryption of the actual UAV location through the public key of the USS. If necessary (see use case #1 in Sec. 3.3), the USS only can decrypt the UAV’s actual location. Note that regular observers do not necessarily need to be updated to be able to decode the new cryptographic fields. General Public Observers can neglect those fields and only refer to the disclosed obfuscated location for the valid drone location. The only observers requiring software updates are the Public Safety Observers, likely connected to the USS for potential disclosure of the drones’ actual location, if necessary. Therefore, our solution preserves backward compatibility with current deployments as much as possible while also enhancing the location privacy of RID-compliant drones.
@@
Utility Considerations. Our experimental assessment (Sec. 8) provides insights into the trade-off between location privacy and utility to LBS. Consider drones obfuscating their location by an average distance of $100$ meters. For the use case #1, CI operators using a small Warning Area achieve above $0.87$ TP rate and only $0.14$ of FP rate. For the use case #2, considering 10 deployed charging stations, the average extra distance travelled by the drone in an area of $1.5$ km ×$\times 2.6$ km ×$\times 40$ m is only $23$ m. Finally, for use case #3, considering the same area and 10 deployed service UAVs, the average extra distance travelled by the signal is only $16$ m. Such results highlight that achieving higher location privacy than the standard RID regulation is possible while also guaranteeing meaningful utility of the disclosed location to LBSs. Recall that, whenever needed, relevant aviation authorities can decrypt the location report to recover the precise drone location.
@@ 10 Conclusion and Future Work
In this paper, we proposed OLO-RID, a framework providing privacy-preserving location disclosure to RID-compliant drones. Through the integration of a state-of-the-art solution for correlated location disclosure and its extension for the drone scenario, OLO-RID allows drones to broadcast through RID messages obfuscated locations, protecting as much as possible trajectory privacy while dealing with processing, communication, memory, and energy constraints of the drones. OLO-RID also extends RID with encrypted location reports, allowing aviation authorities to obtain the actual drone location if necessary. We implemented two proofs-of-concept of our solution on a regular laptop and a RPI, and we experimentally show the viability of our solution for medium-end RID-compliant drones ($0.16$ s of RID message generation time and $0.0236\%$ of energy for a DJI Mini 2). We also evaluate the utility of the disclosed location for three use cases involving drones usage, achieving a reasonable trade-off between drones location privacy and the quality of the provided service. Overall, our work demonstrates that we can achieve accountability for drone operations without entirely sacrificing privacy, paving the way for a practical real-world regulatory framework that considers all such requirements simultaneously.

### v2

@@ 9 Discussion
Summary. Through our extensive performance assessment (Sec. 8), we experimentally proved that OLO-RID allows to generate and deliver RID messages protecting the UAVs’ location privacy within $0.16$ s, i.e., below the threshold of $1$ s set by the RID regulation [remote_id]. The RAM consumption and communication overhead of our solution are also negligible, guaranteeing the potential smooth integration of our solution into running RID-compliant deployments.
[…]
@@
Compliance to Requirements. Here, we look back to the requirements defined in Sec.3.4 to evaluate the compliance of our solution. The requirement R1, Location Obfuscation is fulfilled by integrating and extending the solution in [xiao2015_ccs]. As per the requirement R2, Trajectory Privacy, as shown in Sec. 6, our system is as secure as the PIM method, so it offers trajectory privacy. Our solution does not require any Internet connection on the UAV, so it also fulfils the requirement R3, No Persistent Internet Connection. Internet connection might be required for some use cases discussed in Sec. 3.3, but the absence of such a connection does not prevent the UAV from running OLO-RID. As per the requirement R4, Maximum Message Generation Time., as shown in Sec. 8.2, the maximum time our solution takes to generate a RID message is always less than $0.16$ seconds. Finally, as per the requirement R5, Maximum Communication Overhead, the modified RID message disclosed through OLO-RID is at most 227 bytes, below the MTU of WiFi of $2,304$  bytes [ieee2012].
@@
RID Extension and Backward Compatibility. To protect the UAVs’ location privacy, OLO-RID extends the format of standard RID messages. Instead of delivering the actual UAV location, we deliver an obfuscated location, i.e., a location loosely correlated with the actual one, as described in Sec. 5. We also extend the standard RID message with encrypted location reports, containing the encryption of the actual UAV location through the public key of the USS. If necessary (see use case #1 in Sec. 3.3), the USS only can decrypt the UAV’s actual location. Note that regular observers do not necessarily need to be updated to be able to decode the new cryptographic fields. General Public Observers can neglect those fields and only refer to the disclosed obfuscated location for the valid UAV location. The only observers requiring software updates are the Public Safety Observers, likely connected to the USS for potential disclosure of the UAVs’ actual location, if necessary. Therefore, our solution preserves backward compatibility with current deployments as much as possible while also enhancing the location privacy of RID-compliant UAVs.
@@
Privacy and Utility Considerations. Our experimental assessment (Sec. 8) provides insights into the trade-off between location privacy and utility to LBS. Consider UAVs obfuscating their location by an average distance of $100$ meters. For the use case #1, CI operators using a small Warning Area achieve above $0.87$ TP rate and only $0.14$ of FP rate. For the use case #2, considering 10 deployed charging stations, the average extra distance travelled by the UAV in an area of $1.5$ km $\times 2.6$ km $\times 40$ m is only $23$ m. Finally, for use case #3, considering the same area and 10 deployed service UAVs, the average extra distance travelled by the signal is only $16$ m. Note that such utility metrics are domain-specific, since they show the applicability of our solution to the specific use cases. Overall, such results highlight that achieving higher location privacy than the standard RID regulation is possible while also guaranteeing meaningful utility of the disclosed location to LBSs. Recall that, whenever needed, relevant aviation authorities can decrypt the location report to recover the precise UAV location. We also highlight that, as already demonstrated by several papers in the literature, e.g., [andres2013_ccs] and [mendes2020_pets], the effective location privacy provided by any privacy-preserving mechanism in the presence of frequent broadcasts of location decreases. They also already demonstrate that privacy decreases more when more messsages ($k$) are collected by the adversary. This is a fundamental limitation of the scenario where (location) privacy protection techniques are applied, and not a problem of our solution. To mitigate such a leakage in our scenario, knowing that RID messages are delivered once a second and a static adversary could receive them over a given area, the operator of the UAV could decrease the value of $\epsilon$ on purpose, so that, in the worst case, the effective privacy protection of the $k$ messages disclosed to the adversary does not exceed a given bound.
Limitations. Our solution applies to RPASs, i.e., UAVs with a pilot (either human or CS). For such UAVs, assuming uniform probability of the next movement is realistic, particularly for entertainment applications and video making, where UAVs could move unpredictably in any direction. We acknowledge that, especially for autonomous UAVs on a mission, such an assumption could be less realistic, requiring a change in the values of the cells in the transition matrix. Although such values do not change the overhead of our solution, they could affect the privacy provided to the autonomous UAV. The performance evaluation presented in Sec. 8.2 uses a single NATO dataset, available at [flight_data], since it is the only publicly available dataset of real UAV flights at the time of this writing. To the best of our knowledge, all other UAV flight datasets available online include synthetic trajectories, which do not represent actual UAV flights. More complex flight paths could change the reported performance and utility figures. We also acknowledge that our paper considers only three reference use cases, and it does not cover all possible use cases where UAVs could be applied.
@@ 10 Conclusion and Future Work
In this paper, we proposed OLO-RID, a framework providing privacy-preserving location disclosure to RID-compliant UAVs. Through the integration of a state-of-the-art solution for correlated location disclosure and its extension for the UAV scenario, OLO-RID allows UAVs to broadcast through RID messages obfuscated locations, protecting as much as possible trajectory privacy while dealing with processing, communication, memory, and energy constraints of the UAVs. OLO-RID also extends RID with encrypted location reports, allowing aviation authorities to obtain the actual UAV location if necessary. We implemented two proofs-of-concept of our solution on a regular laptop and a RPI, and we experimentally show the viability of our solution for medium-end RID-compliant UAVs ($0.16$ s of RID message generation time and $0.0236\%$ of energy for a DJI Mini 2). We also evaluate the utility of the disclosed location for three use cases involving UAVs usage, achieving a reasonable trade-off between UAVs location privacy and the quality of the provided service. Overall, our work demonstrates that we can achieve accountability for UAV operations without entirely sacrificing privacy, paving the way for a practical real-world regulatory framework that considers all such requirements simultaneously.